# The appearance of the regular black hole with a stable Cauchy horizon

### 07 Dec 2023 v1

@@ Abstract
The strong cosmic censorship conjecture, which states that the evolution of generic initial data will always produce a globally hyperbolic spacetime, is hard to be tested by astronomical observations. In this paper, we study the appearance of the regular black hole without mass inflation, which violates the strong cosmic censorship conjecture. Since the Cauchy horizon is stable, the photons entering the two horizons of the regular black hole in the preceding companion universe can come out from the white hole in our universe. These rays create a novel multi-ring structure, which is significantly different from the image of the Schwarzschild black hole. This serves a potential method to test the strong cosmic censorship conjecture.
@@ I Introduction
It has been years since the images of the supermassive objects at the center of M87 EventHorizonTelescope:2019dse and Milky Way galaxies EventHorizonTelescope:2022wkp was captured by the Event Horizon Telescope (EHT). These provide the evidence for the existence of black holes. In these images, the dark area is surrounded by a bright ring-like structure, which is illuminated dominantly by the direct emission of the accretion disk. In theoretical calculations, there is a narrow bright ring besides the direct emission from the accretion disk Gralla:2019xty . This is so called lensed ring, which is constituted by the ray deflected at least half an orbit around the object. When the impact parameter approaches the critical value, the photons will rotate several times around the black hole before creating a curve on the observer’s image plane. As the impact parameter reaches its critical limit, the number of rotations increases infinitely, causing the light ray to approach the critical curve. Consequently, these rays form the photon ring. However, it is so close to the lensed ring that it can not be distinguished, and its contribution to the overall intensity is negligible. Although the critical curve and the photon ring are fully determined by the background geometry, the optical appearance of a compact object is heavily influenced by the position and profile of light source. This gives us a window through which we can get the background geometry and accretion disk information from the black hole image.
@@
Last century, Hawking and Penrose demonstrated that singularities occur unavoidably in general relativity (GR) Penrose:1964wq ; Hawking:1970zqf . Weak cosmic censorship conjecture (WCCC) Penrose:1969pc postulate that singularities, where physical laws break down, always hide behind the horizon. Furthermore, the presence of the singularity, i.e., the existence of incomplete geodesics, threatens the predictability of GR. Therefore, regular black holes without singularities, which are constructed by relaxing some conditions in the singularity theorem, attract much attention. Following Bardeen’s work Bardeen , many regular black holes were discovered, such as the loop black hole Modesto:2008im , the Hayward black hole Hayward:2005gi , the quantum-corrected black hole Lewandowski:2022zce and so on Bambi:2023try ; Lan:2023cvz . Regular black holes are usually constructed by replacing the central singularity with a non-singular core. They are classified as de Sitter core and Minkowski core according to the asymptotic behavior near $r=0$. Even though regular black holes avoid singularities at their centers, most of them possess an unstable Cauchy horizon. In fact, all spherically symmetric regular black holes have inner horizons Carballo-Rubio:2019fnb . If the spacetime can extend beyond the Cauchy horizon, then it allows the infinite repetitions within the Penrose diagram of the spacetime. For example, the Penrose diagram of the Reissner-Nordström (RN) black hole is consisted of countless repeated identical universes. This will break down the predictability of GR, and the classical determinism is lost. Fortunately, most of these inner horizons of regular black holes are unstable, which are caused by mass inflation and will convert into null singularities Brown:2011tv ; Iofa:2022dnc ; Carballo-Rubio:2021bpr . The instability of the inner horizon prevents the spacetime extending beyond it, and thus avoid the problem of predictability. This picture supports the strong cosmic censorship conjecture (SCCC), which states that a physical spactime is always globally hyperbolic Penrose:1978 ; Penrose:1979azm . In the modern language of partial differential equations, the evolution of generic initial data will always produce a globally hyperbolic spacetime. In other words, a stable Cauchy horizon is in contradiction with the SCCC. Examination of the SCCC is an important topic in mathematical physics nowadays Dafermos:2002ka ; Dafermos:2003vim ; Dafermos:2003wr ; Dafermos:2012np ; Dafermos:2017dbw . However, it is challenging to verify this conjecture through astronomical observations due to the lack of proper ways.
@@
The image of the regular black hole has been studied in recent years Guo:2021wid ; DeMartino:2023ovj . Their shapes are similar to the case of the Schwarzschild black hole, except for some differences in intensity, position of the photon ring and so on. Because the images of the regular black holes in Guo:2021wid ; DeMartino:2023ovj and the Schwarzschild black hole are mainly determined by the direct emission, which is strongly influenced by the accretion disk, it is not easy to distinguish the images of regular black holes with the Schwarzschild one. It is exciting that the quantum-corrected black hole may have a novel image Zhang:2023okw . This is a regular black hole proposed by the quantum Oppenheimer-Snyder and Swiss Cheese models in the framework of Loop Quantum Gravity (LQG) Lewandowski:2022zce . The photon can enter the event horizon of a companion black hole in the preceding universe, come out from the white hole in our universe and finally be captured by the observer. These rays produces many new bright rings in the image of the black hole, especially some of these rings are distinctly visible within the shadow. This multi-ring structure may be detected astronomically. However, it has been proved that the Cauchy horizon of the quantum-corrected black hole is unstable in Cao:2023aco . After a small perturbation, the mass inflation occurs and the Cauchy horizon becomes a null singularity. This prevents the spacetime to extend beyond the Cauchy horizon, and the photons can not pass though it. Therefore, whether this novel multi-ring structure will occur is still unclear.
@@
Intriguingly, a regular black hole with a stable Cauchy horizon was proposed in Carballo-Rubio:2022kad . The vanishing surface gravity of the Cauchy horizon eliminates the possibility of the mass inflation instability. Using the ray-tracing method Gralla:2019xty , we draw its image in section III. Similar to the quantum-corrected black hole, which ignoring the instability of the Cauchy horizon, a multi-ring structure emerges. The stable Cauchy horizon ensures that the photons can pass though the two horizons in the companion black hole in the preceding universe, and come out from the white hole in our universe. Therefore, unlike the quantum-corrected black hole, this multi-ring structure is physically permissible. However, as we mentioned above, the stable Cauchy horizon will violate the SCCC. Hence, the presences of multi-ring structure in observation may suggest that SCCC is violated. This provides us a way to test the SCCC in astronomical observations.
@@
This paper is organized as follows. In section II, we introduce the regular black hole with a stable Cauchy horizon and get its Penrose diagram. We then consider thin disk emission at different locations and investigate their observational appearance in section III. In section IV, the image of the regular black with different parameters is investigated. Finally, we give conclusions and discussions in section V.
II The regular black hole with a stable Cauchy horizon
A regular black hole with a stable Cauchy horizon is proposed in Carballo-Rubio:2022kad . In Schwarzschild-like coordinates, the metric of this spherically symmetric regular black hole can be written as
@@ II The regular black hole with a stable Cauchy horizon
$M$ is the mass parameter of the black hole, $r_{+}$ and $r_{-}$ are the radii of the event horizon and the Cauchy horizon of the black hole respectively. The parameter $b_{2}$ is sufficiently large and is related to the parameter $a_{2}$ as follows
@@
Figure 1: The metric function $f(r)$ of the regular black hole (solid curve) with $M=1,r_{-}=0.1,r_{+}=2,b_{2}=5$ and the Schwarzschild black hole (dashed curve) with $M=1$.
@@
$f\rightarrow 1-\frac{b_{2}}{r_{-}^{2}r_{+}}r^{2}\,.$ (4)
It is worth noting that the surface gravity of the Cauchy horizon vanishes, i.e., $\kappa_{-}=0$, while that of the event horizon is nonzero, i.e., $\kappa_{+}\neq 0$. Therefore, the observer will receive a finite flux near the Cauchy horizon after a small perturbation Chandrasekhar and Hartle (1982). Besides, based on the model by Ori Ori:1991zz and the general DTR relation Brown:2011tv , it can be shown that the Cauchy horizon is stable Carballo-Rubio:2022kad .
@@
Figure 2: The block with $r<r_{-}$ of the RN black hole.
@@
Figure 3: The Penrose diagram of the regular black hole. $r=0$ is not a singularity.
@@
Figure 4: The effective potential $V(r)$ of the regular black hole (solid curve) with $M=1,r_{-}=0.1,r_{+}=2,b_{2}=5$ and the Schwarzschild black hole (dashed curve) with $M=1$.
@@ III The ray-tracing method and the appearance of the regular black hole
Figure 5: The photon trajectory in the Penrose diagram. The two dashed curve in universe $A$ and $B$ is the world lines of the critical curve. The gray curve is the photon trajectory from the critical curve to the observer in the same universe $B$. And the black curve is the photon trajectory from the critical curve in $A$ to the observer in $B$.
The accretion disk is around the black hole (or white hole) and is sharply peaked at the critical curve, which is also called “photon sphere” where the light orbits around it infinite times. In Fig.5 the critical curve or photon sphere is described by the dashed curve. As we can see, the observer in universe $B$ can receive the photons from two sources. One is from the accretion disk in universe $B$, illustrated as the gray curve in Fig.5. Another is showed as the black curve in Fig.5. The photons are emitted from the the accretion disk in universe $A$, subsequently falling into the event horizon, then traverse the Cauchy horizon, and arrive at the universe $D$. Then they travel through the Cauchy horizon and the event horizon of the white hole in turn, and are finally captured by the observer in universe $B$. These two kinds of source of photons lead to different images. Therefore, we will discuss these two cases separately.
@@
We assume the light emitted is monochromatic, so the photons are emitted with the same frequency. The emission is sharply peaked and ends at the critical curve, as shown in Fig.6
Figure 6: The emission profiles with $r_{m}=3.014$. Here $r_{m}$ is the critical curve of the regular black hole with $r_{+}=2,r_{-}=1/50,M=1,b_{2}=0$.
@@ III.1 The image of the black hole formed by the photons emitted in universe $B$
Figure 7: The trajectory of the photon around the regular black hole with $r_{+}=2,r_{-}=1/50,M=1,b_{2}=0$. The green line denotes the observation screen at the null infinity. The black circle corresponds to the critical curve and the blue circle corresponds to the event horizon. The dashed line is the accretion disk, which is the light source and located at the equatorial plane. The colored curves correspond to $n<3/4$ for red, 3/$43/4<n<5/$ 4 for orange, $n>5/4$ for yellow, and $b<b_{c}$ for black.
The normalized number of orbits $n=\phi/(2\pi)$ relates to the number of intersections with the equatorial plane of a particular light ray, where $\phi$ is the azimuthal angle. The null geodesic with $b>6.215$ is shown as red curves with $n<3/4$. They intersect with the equatorial plane (dashed line) for one time, so is called “direct emission”. The null geodesic with $6.2155.264<b<$ 6.215 is colored by orange with 3/$43/4<n<5/$ 4. They intersect with the equatorial plane for two times and are called “lensed ring”. At each intersection, newly emitted photons from the accretion disk join in the journey towards the screen. The null geodesic with $5.2645.230<b<$ 5.264 is colored by yellow with $n>5/4$. They cross the equatorial plane more than 3 times and is called “photon ring”. The impact parameter of the critical curve is $b_{c}=5.231$. The rays with $b<b_{c}$ is direct emission, which is similar to the situation of $b>6.215$ and is shown as the black curves. And the rays with $b<3.904$ can not intersect with the accretion disk and, therefore, do not contribute to the intensity.
@@ III.2 The image of the black hole formed by the photons emitted in universe $A$
Since the Cauchy horizon is stable, the photons falling into the event horizon can cross the Cauchy horizon. Therefore, when the accretion disk is located at universe $A$, the photon with $b<b_{c}$ can also be received by the observer in universe $B$, as the black curve in Fig.5. In this case, the trajectory of the photon outside the white hole in universe $B$ is shown in Fig.9.
@@
The inner circle and the outer circle are the event horizon and the critical curve of the white hole in universe $B$, respectively.
@@
Figure 10: The trajectory of the photons from universe $A$ to universe $B$. The black circle and the blue circle are the critical curve and the event horizon of the black (white) hole.
@@
Figure 11: The trajectory of the photons with some critical $n$. All the ray end at $x\rightarrow+\infty$. Picture (b) shows the stretching of the $x$-axis in picture (a). And picture (c) shows the stretching of the $y$-axis in picture (b). The black circle is the critical curve and the blue circle is the event horizon. The red curve and purple curve correspond to $n=3/4$ and $5/4$. The orange curve intersect with the accretion disk and the critical curve at the same point C. And the green curve intersect with them at D.
@@
Therefore, as $b$ increases, the azimuthal angle increase, too. When $b=0.02454$, the azimuthal angle $\phi$ of the starting point of the ray is $(3/2)\pi$, or $n=3/4$, as the red curve. If we increase $b$ slightly, then the ray will have a larger deflection angle, and intersects with the accretion disk. This kind of ray can be received by the screen, until $b=0.02461$ as the orange curve. Therefore, a bright ring occurs on the screen. However, as $b$ continues to increase, the ray will not intersect with the accretion disk. Until the ray intersects with the accretion disk again at the top half of the $y$-axis with $n\gtrsim 5/4,b\gtrsim 0.05158$ as the purple curve, and the second bright ring occurs. This second ring end at $b=0.05174$ as the green curve, which intersect with the accretion disk at point $D$. In the same manner, there are many rings appear on the screen with $b<b_{c}$. As the value of $r$ approaches zero, $V(r)$ tends to $1/r^{2}$. Consequently, the trajectory of the light ray we mentioned above behaves similar to that of Cotes’s spirals Whittaker:1988 . Although there being only a single turning point, the light ray has the capability to rotate around the center multiple times.
@@
As we can see, the multi-ring structure occurs for the regular black hole. Unlike the image of the Schwarzschild black hole, where there is only one or two bright rings lying in the region $b\gtrsim b_{c}$ when the accretion disk located outside the critical curve, in our case, the bright rings with $n>5/4$ and $b<b_{c}$ are distinctly separated, allowing us to distinguish them. When $b$ tends to $b_{c}$, the bright rings become more and more concentrated and indistinguishable. This multi-ring structure also appears in some other massive bodies without horizon, such as compact object and wormhole Olmo:2021piq ; Guerrero:2022qkh ; Guerrero:2022msp ; Olmo:2023lil . In fact, when the effective potential $V(r)$ of the spacetime have a barrier, the critical curve occurs and there is a bright ring in the image. Furthermore, if there is an another higher barrier inside the outer one, or $V$ is diverging at somewhere inside the critical curve, then the multi-ring structure appears. In Zhang:2023okw , the image of the quantum-corrected black hole is investigated, and it has the similar multi-ring structure. However, the Cauchy horizon of the quantum-corrected black hole is unstable and will turn into a null singularity Cao:2023aco . Therefore the photons are unable to pass through the inner horizon and escape from a white hole. Due to the instability of its Cauchy horizon and the slightness of the quantum correction, it may be challenging to detect this novel appearance of the quantum-corrected black hole. Nevertheless, the Cauchy horizon of the regular black hole we are studying is stable which give rise to the multi-ring structure. It also means that the multi-ring structure can occur in the appearance of the massive object with horizon.

### 23 Feb 2024 v2

# The appearance of the regular black hole with a stable inner horizon

@@ Abstract
The strong cosmic censorship conjecture, which states that the evolution of generic initial data will always produce a globally hyperbolic spacetime, is hard to be tested by astronomical observations. In this paper, we study the appearance of the regular black hole without mass inflation, which violates the strong cosmic censorship conjecture. Since the inner horizon is stable, the photons entering the two horizons of the regular black hole in the preceding companion universe can come out from the white hole in our universe. These rays create a novel multi-ring structure, which is significantly different from the image of the Schwarzschild black hole. This serves a potential method to test the strong cosmic censorship conjecture.
@@ I Introduction
It has been years since the images of the supermassive objects at the center of M87 EventHorizonTelescope:2019dse and Milky Way galaxies EventHorizonTelescope:2022wkp was captured by the Event Horizon Telescope (EHT). These provide the evidence for the existence of black holes. In these images, the dark area is surrounded by a bright disk-like structure, which is illuminated dominantly by the direct emission of the accretion disk. In theoretical calculations, there is a narrow bright ring besides the direct emission from the accretion disk Gralla:2019xty . This is so called lensing ring, which is constituted by the ray emitted near the photon sphere. The photon sphere is a region of spacetime where photons can orbit a black hole an infinite number of times. In the image plane, the curve corresponding to the photon sphere is referred to as the critical curve. Surrounding this critical curve in the image plane, there is a region of enhanced brightness. Within this region, two distinct phenomena occur: the photon ring and the lensing ring. The photon ring is formed by light rays that intersect the accretion disk more than twice, while the lensing ring is formed by light rays that intersect the accretion disk exactly twice. However, the photon ring is so close to the lensing ring that it can not be distinguished, and its contribution to the overall intensity is negligible. Although the critical curve and the photon ring are fully determined by the background geometry, the optical appearance of a compact object is heavily influenced by the position and profile of light source. This gives us a window through which we can get the background geometry and accretion disk information from the black hole image.
@@
Last century, Hawking and Penrose demonstrated that singularities occur unavoidably in general relativity (GR) Penrose:1964wq ; Hawking:1970zqf . Weak cosmic censorship conjecture (WCCC) Penrose:1969pc postulate that singularities, where physical laws break down, always hide behind the horizon. Furthermore, the presence of the singularity, i.e., the existence of incomplete geodesics, threatens the predictability of GR. Therefore, regular black holes without singularities, which are constructed by relaxing some conditions in the singularity theorem, attract much attention. Following Bardeen’s work Bardeen , many regular black holes were discovered, such as the loop black hole Modesto:2008im , the Hayward black hole Hayward:2005gi , the quantum-corrected black hole Lewandowski:2022zce and so on Bambi:2023try ; Lan:2023cvz . Regular black holes are usually constructed by replacing the central singularity with a non-singular core. They are classified as de Sitter core and Minkowski core according to the asymptotic behavior near $r=0$. The WCCC is discussed in the rotate regular black hole in Ghosh:2022gka . Even though regular black holes avoid singularities at their centers, most of them possess an unstable inner horizon. In fact, all spherically symmetric regular black holes have inner horizons Carballo-Rubio:2019fnb . If the Cauchy horizon is stable, the Penrose diagram of the spacetime can be extended to consist of infinite repeated universes, thereby implying an infinite series of discrete inner horizons. The Cauchy horizon of a certain initial data surface is just a part of certain inner horizon. If we confine ourselves to a single universe, the inner horizon can be identical to the Cauchy horizon. Hereafter, we will mainly use the terminology “inner horizon” instead of “Cauchy horizon” since their subtle distinctions won’t be crucial for our discussion. The extension beyond Cauchy horizon will break down the predictability of GR, and the classical determinism is lost. Fortunately, most of these inner horizons of regular black holes are unstable, which are caused by mass inflation and will convert into null singularities Brown:2011tv ; Iofa:2022dnc ; Carballo-Rubio:2021bpr . The instability of the inner horizon prevents the spacetime extending beyond it, and thus avoid the problem of predictability. This picture supports the strong cosmic censorship conjecture (SCCC), which states that a physical spactime is always globally hyperbolic Penrose:1978 ; Penrose:1979azm . In the modern language of partial differential equations, the evolution of generic initial data will always produce a globally hyperbolic spacetime. In other words, a stable Cauchy horizon is in contradiction with the SCCC. Examination of the SCCC is an important topic in mathematical physics nowadays Dafermos:2002ka ; Dafermos:2003vim ; Dafermos:2003wr ; Dafermos:2012np ; Dafermos:2017dbw . However, it is challenging to verify this conjecture through astronomical observations due to the lack of proper ways.
@@
The image of the regular black hole has been studied in recent years Guo:2021wid ; DeMartino:2023ovj ; Tsukamoto:2017fxq ; Tsukamoto:2014tja ; Jiang:2023img ; Ghosh:2022gka . Their shapes are similar to the case of the Schwarzschild black hole or the Kerr black hole, except for some differences in intensity, position of the photon ring and so on. Because the images of the regular black holes in Guo:2021wid ; DeMartino:2023ovj and the Schwarzschild black hole are mainly determined by the direct emission, which is strongly influenced by the accretion disk, it is not easy to distinguish the images of regular black holes with the Schwarzschild one. It is exciting that the quantum-corrected black hole may have a novel image Zhang:2023okw . This is a regular black hole proposed by the quantum Oppenheimer-Snyder and Swiss Cheese models in the framework of Loop Quantum Gravity (LQG) Lewandowski:2022zce . The photon can enter the event horizon of a companion black hole in the preceding universe, come out from the white hole in our universe and finally be captured by the observer. These rays produces many new bright rings in the image of the black hole, especially some of these rings are distinctly visible within the shadow. This multi-ring structure may be detected astronomically. However, it has been proved that the inner horizon of the quantum-corrected black hole is unstable in Cao:2023aco . After a small perturbation, the mass inflation occurs and the inner horizon becomes a null singularity. This prevents the spacetime to extend beyond the inner horizon, and the photons can not pass though it. Therefore, whether this novel multi-ring structure will occur is still unclear.
@@
Intriguingly, a regular black hole with a stable inner horizon was proposed in Carballo-Rubio:2022kad . The vanishing surface gravity of the inner horizon eliminates the possibility of the mass inflation instability. Using the ray-tracing method Gralla:2019xty , we draw its image in section III. Similar to the quantum-corrected black hole, which ignoring the instability of the inner horizon, a multi-ring structure emerges. The stable inner horizon ensures that the photons can pass though the two horizons in the companion black hole in the preceding universe, and come out from the white hole in our universe. Therefore, unlike the quantum-corrected black hole, this multi-ring structure is physically permissible. However, as we mentioned above, the stable Cauchy horizon will violate the SCCC. Hence, the presences of multi-ring structure in observation may suggest that SCCC is violated. This provides us a way to test the SCCC in astronomical observations.
@@
This paper is organized as follows. In section II, we introduce the regular black hole with a stable inner horizon and get its Penrose diagram. We then consider thin disk emission at different locations and investigate their observational appearance in section III. In section IV, the image of the regular black with different parameters is investigated. Finally, we give conclusions and discussions in section V.
II The regular black hole with a stable inner horizon
A regular black hole with a stable inner horizon is proposed in Carballo-Rubio:2022kad . In Schwarzschild-like coordinates, the metric of this spherically symmetric regular black hole can be written as
@@ II The regular black hole with a stable inner horizon
$M$ is the mass parameter of the black hole, $r_{+}$ and $r_{-}$ are the radii of the event horizon and the inner horizon of the black hole respectively. The parameter $b_{2}$ is sufficiently large and is related to the parameter $a_{2}$ as follows
@@
Figure 1: The metric function $f(r)$ (a) and the effective potential (b) of the regular black hole (solid curve) with $M=1,r_{-}=0.1,r_{+}=2,b_{2}=5$ and the Schwarzschild black hole (dashed curve) with $M=1$.
@@
$f\rightarrow 1-\frac{b_{2}}{r_{-}^{3}r_{+}}r^{2}\,.$ (4)
It is worth noting that the surface gravity of the inner horizon vanishes, i.e., $\kappa_{-}=0$, while that of the event horizon is nonzero, i.e., $\kappa_{+}\neq 0$. Therefore, the observer will receive a finite flux near the inner horizon after a small perturbation Chandrasekhar and Hartle (1982). Besides, based on the model by Ori Ori:1991zz and the general DTR relation Brown:2011tv , it can be shown that the inner horizon is stable Carballo-Rubio:2022kad .
@@
Figure 2: (a) is the block with $r<r_{-}$ of the RN black hole and (b) is the Penrose diagram of the regular black hole. $r=0$ of the regular black hole is not a singularity.
@@ III The ray-tracing method and the appearance of the regular black hole
Figure 3: The photon trajectory in the Penrose diagram. The two dashed curve in universe $A$ and $B$ is the world lines of the photon sphere. The gray curve is the photon trajectory from the photon sphere to the observer in the same universe $B$. And the black curve is the photon trajectory from the photon sphere in $A$ to the observer in $B$.
The accretion disk is around the black hole (or white hole) and is sharply peaked at the photon sphere, which is also called “photon sphere” where the light orbits around it infinite times. In Fig.3 the photon sphere is described by the dashed curve. As we can see, the observer in universe $B$ can receive the photons from two sources. One is from the accretion disk in universe $B$, illustrated as the gray curve in Fig.3. Another is showed as the black curve in Fig.3. The photons are emitted from the the accretion disk in universe $A$, subsequently falling into the event horizon, then traverse the inner horizon, and arrive at the universe $D$. Then they travel through the inner horizon and the event horizon of the white hole in turn, and are finally captured by the observer in universe $B$. These two kinds of source of photons lead to different images. Therefore, we will discuss these two cases separately.
@@
We assume the light emitted is monochromatic, so the photons are emitted with the same frequency. The emission is sharply peaked and ends at the photon sphere, as shown in Fig.4
Figure 4: The emission profiles with $r_{m}=3.014$. Here $r_{m}$ is the photon sphere of the regular black hole with $r_{+}=2,r_{-}=1/50,M=1,b_{2}=0$.
@@ III.1 The image of the black hole formed by the photons emitted in universe $B$
Figure 5: The trajectory of the photon around the regular black hole with $r_{+}=2,r_{-}=1/50,M=1,b_{2}=0$. The green line denotes the observation screen at the null infinity. The black circle corresponds to the photon sphere and the blue circle corresponds to the event horizon. The dashed line is the accretion disk, which is the light source and located at the equatorial plane. The colored curves correspond to $n<3/4$ for red, 3/$43/4<n<5/$ 4 for orange, $n>5/4$ for yellow, and $b<b_{c}$ for black.
The normalized number of orbits $n=\phi/(2\pi)$ relates to the number of intersections with the equatorial plane of a particular light ray, where $\phi$ is the azimuthal angle. The null geodesic with $b>6.215$ is shown as red curves with $n<3/4$. They intersect with the equatorial plane (dashed line) for one time, so is called “direct emission”. The null geodesic with $6.2155.264<b<$ 6.215 is colored by orange with 3/$43/4<n<5/$ 4. They intersect with the equatorial plane for two times and are called “lensing ring”. At each intersection, newly emitted photons from the accretion disk join in the journey towards the screen. The null geodesic with $5.2645.230<b<$ 5.264 is colored by yellow with $n>5/4$. They cross the equatorial plane more than 3 times and is called “photon ring”. The impact parameter of the critical curve is $b_{c}=5.231$. The rays with $b<b_{c}$ is direct emission, which is similar to the situation of $b>6.215$ and is shown as the black curves. And the rays with $b<3.904$ can not intersect with the accretion disk and, therefore, do not contribute to the intensity.
@@ III.2 The image of the black hole formed by the photons emitted in universe $A$
Since the inner horizon is stable, the photons falling into the event horizon can cross the inner horizon. Therefore, when the accretion disk is located at universe $A$, the photon with $b<b_{c}$ can also be received by the observer in universe $B$, as the black curve in Fig.3. In this case, the trajectory of the photon outside the white hole in universe $B$ is shown in Fig.7.
@@
The inner circle and the outer circle are the event horizon and the photon sphere of the white hole in universe $B$, respectively.
@@
Figure 8: The trajectory of the photons from universe $A$ to universe $B$. The black circle and the blue circle are the photon sphere and the event horizon of the black (white) hole.
@@
Figure 9: The trajectory of the photons with some critical $n$. All the ray end at $x\rightarrow+\infty$. Picture (b) shows the stretching of the $x$-axis in picture (a). And picture (c) shows the stretching of the $y$-axis in picture (b). The black circle is the photon sphere and the blue circle is the event horizon. The red curve and purple curve correspond to $n=3/4$ and $5/4$. The orange curve intersect with the accretion disk and the photon sphere at the same point $D$. And the green curve intersect with them at $C$.
@@
Therefore, as $b$ increases, the azimuthal angle increase, too. When $b=0.02454$, the azimuthal angle $\phi$ of the starting point of the ray is $(3/2)\pi$, or $n=3/4$, as the red curve. If we increase $b$ slightly, then the ray will have a larger deflection angle, and intersects with the accretion disk. This kind of ray can be received by the screen, until $b=0.02461$ as the orange curve. Therefore, a bright ring occurs on the screen. However, as $b$ continues to increase, the ray will not intersect with the accretion disk. Until the ray intersects with the accretion disk again at the top half of the $y$-axis with $n\gtrsim 5/4,b\gtrsim 0.05158$ as the purple curve, and the second bright ring occurs. This second ring end at $b=0.05174$ as the green curve, which intersect with the accretion disk at point $C$. In the same manner, there are many rings appear on the screen with $b<b_{c}$. As the value of $r$ approaches zero, $V(r)$ tends to $1/r^{2}$. Consequently, the trajectory of the light ray we mentioned above behaves similar to that of Cotes’s spirals Whittaker:1988 . Although there being only a single turning point, the light ray has the capability to rotate around the center multiple times.
@@
As we can see, the multi-ring structure occurs for the regular black hole. Unlike the image of the Schwarzschild black hole, where there is only one or two bright rings lying in the region $b\gtrsim b_{c}$ when the accretion disk located outside the photon sphere, in our case, the bright rings with $n>5/4$ and $b<b_{c}$ are distinctly separated, allowing us to distinguish them. When $b$ tends to $b_{c}$, the bright rings become more and more concentrated and indistinguishable. This multi-ring structure also appears in some other massive bodies without horizon, such as compact object and wormhole Olmo:2021piq ; Guerrero:2022qkh ; Guerrero:2022msp ; Olmo:2023lil . In fact, when the effective potential $V(r)$ of the spacetime have a barrier, the photon sphere occurs and there is a bright ring in the image. Furthermore, if there is an another higher barrier inside the outer one, or $V$ is diverging at somewhere inside the photon sphere, then the multi-ring structure appears. In Zhang:2023okw , the image of the quantum-corrected black hole is investigated, and it has the similar multi-ring structure. However, the inner horizon of the quantum-corrected black hole is unstable and will turn into a null singularity Cao:2023aco . Therefore the photons are unable to pass through the inner horizon and escape from a white hole. Due to the instability of its inner horizon and the slightness of the quantum correction, it may be challenging to detect this novel appearance of the quantum-corrected black hole. Nevertheless, the inner horizon of the regular black hole we are studying is stable which give rise to the multi-ring structure. It also means that the multi-ring structure can occur in the appearance of the massive object with horizon.